# Flow-based Conformal Prediction for Multi-dimensional Time Series

### v2

@@ 6 Conclusion
In this study, we proposed a novel conformal prediction method for multi-dimensional time series using flow with classifier-free guidance. We provided coverage guarantees of our method by establishing exact non-asymptotic marginal coverage and a finite-sample bound on conditional coverage. Experiments on real-world datasets with a broad set of baselines demonstrated that our method constructs smaller prediction sets while satisfying the target coverage, consistently outperforming the baselines. Future work will investigate dynamic optimal transport mappings, implemented through flow, between the non-conformity scores and the source distribution, with the aim of constructing more efficient prediction sets and deriving sharper coverage bounds.
@@ Appendix A Proofs
Since $u_{t|x_{1}}(x\mid x_{1})$ generates the probability path $p_{t|x_{1}}(x\mid x_{1})$, the continuity equation holds for each $x_{1}$:
$\frac{\partial p_{t|x_{1}}(x\mid x_{1})}{\partial t}+\text{div}\left(u_{t|x_{1}}(x\mid x_{1})p_{t|x_{1}}(x\mid x_{1})\right)=0.$ (31)
The time derivative of $p_{t|h}(x\mid h)$ is:
$\frac{\partial p_{t|h}(x\mid h)}{\partial t}$ $=\frac{\partial}{\partial t}\int p_{t|x_{1}}(x\mid x_{1})q(x_{1}\mid h)\,dx_{1}$ (32)
$=\int\frac{\partial p_{t|x_{1}}(x\mid x_{1})}{\partial t}q(x_{1}\mid h)\,dx_{1}$
@@
$=-\text{div}\left(\int u_{t|x_{1}}(x\mid x_{1})p_{t|x_{1}}(x\mid x_{1})q(x_{1}\mid h)\,dx_{1}\right).$
Since the marginal guided vector field is defined as:
$u_{t|h}(x\mid h):=\int u_{t|x_{1}}(x\mid x_{1})\frac{p_{t|x_{1}}(x\mid x_{1})q(x_{1}\mid h)}{p_{t|h}(x\mid h)}\,dx_{1},$ (33)
we can rewrite as:
@@
$\frac{\partial p_{t|h}(x\mid h)}{\partial t}=-\text{div}\left(u_{t|h}(x\mid h)p_{t|h}(x\mid h)\right),$ (35)
which is the continuity equation for $p_{t|h}(x\mid h)$ under the vector field $u_{t|h}(x\mid h)$. Therefore, $u_{t|h}(x\mid h)$ is a valid vector field generating $p_{t|h}(x\mid h)$. ∎
@@
By the definition of the guided marginal probability path:
@@
where $p_{t|x_{1}}(x\mid x_{1})=\mathcal{N}(x\mid\alpha_{t}x_{1},\sigma_{t}^{2}I)$. We express the score function as
@@
$=\frac{\int\nabla_{x}p_{t|x_{1}}(x\mid x_{1})q(x_{1}\mid y)dx_{1}}{p_{t|h}(x\mid h)}$ (39)
$=\int\nabla_{x}\log p_{t|x_{1}}(x\mid x_{1})\frac{p_{t|x_{1}}(x\mid x_{1})q(x_{1}\mid y)}{p_{t|h}(x\mid h)}dx_{1}.$ (40)
@@
$u_{t}(x\mid x_{1})$ $=\frac{\dot{\alpha}_{t}}{\sigma_{t}}(x-\alpha_{t}x_{1})+\dot{\alpha}_{t}x_{1}$ (41)
@@
$=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+(\dot{\alpha}_{t}\sigma_{t}-\alpha_{t}\dot{\sigma}_{t})\frac{\sigma_{t}}{\alpha_{t}}\nabla\log p_{t}(x\mid x_{1}),$ (45)
@@
Therefore,
@@
By using the identity $\nabla_{x}\log p_{t|h}(x\mid h)=\nabla_{x}\log p_{h|t}(h\mid x)+\nabla_{x}\log p_{t}(x)$, we have:
@@
By using Jacobi’s formula,
@@
$\text{tr}\left(\frac{\partial u_{t|h}(\psi_{t|h}(x\mid h))}{\partial\psi_{t|h}(x\mid h)}\right)=\text{div}\left(u_{t|h}(\psi_{t|h}(x\mid h))\right),$ (54)
where $\text{div}(\cdot)$ denotes the divergence operator.
@@
Theorem A.4 (Closed and connected sets under a continuous map, munkres2000topology).
Let $Z$ and $Y$ be topological spaces, and let $\psi:Z\to Y$ be a continuous map. If $E\subset Z$ is closed and connected, then $\psi(E)\subset Y$ is also closed and connected.
Assumption A.5 (Compact feature and outcome domains).
The feature and outcome domains are compact. That is, $x_{i}\in\mathcal{X}\subset\mathbb{R}^{d_{x}}$ and $y_{i}\in\mathcal{Y}\subset\mathbb{R}^{d_{y}}$, where $\mathcal{X}$ and $\mathcal{Y}$ are compact sets.
Remark A.6.
While its not strictly required, further assume that the domains of $x_{i}$ and $y_{i}$ are compact, which ensures that the encoder output is also compact, as formalized in Assumption A.5. Under Assumption A.5, if the encoder is a continuous function that maps a sequence of inputs to a representation $h\in\mathbb{R}^{d_{h}}$, then the image of the encoder $\mathcal{H}\subset\mathbb{R}^{d_{h}}$ is compact.
@@
Let $\psi_{t}$ denote the guided flow defined by a guided vector field $u_{t}$. If the guided vector field $u_{t}(x\mid h)$ is Lipschitz continuous in $x$ uniformly over $t\in[0,1]$ and $h\in\mathcal{H}$, i.e., there exists a constant $L_{u}>0$ such that
@@
then the guided flow $\psi_{t}(x\mid h)$ is Lipschitz continuous in $x$ over $t\in[0,1]$ and $h\in\mathcal{H}$. That is, there exists a constant $L_{\psi}>0$ such that
@@
Let $d(t)=\|\psi_{t}(x\mid h)-\psi_{t}(x^{\prime}\mid h)\|$
Since the guided vector field is Lipschitz continuous, there exists $L_{u}$ such that
@@
Let $z(t)=\psi_{t}(x\mid h)-\psi_{t}(x^{\prime}\mid h)$, then
$\frac{d}{dt}d(t)=\frac{1}{\|z(t)\|}\langle z(t),\frac{d}{dt}z(t)\rangle=\langle\frac{z(t)}{\|z(t)\|},\frac{d}{dt}z(t)\rangle$ (61)
Since $\frac{d}{dt}z(t)=u_{t}(\psi_{t}(x\mid h)\mid h)-u_{t}(\psi_{t}(x^{\prime}\mid h)\mid h)$, by Cauchy-Schwarz inequality,
$|\langle\frac{z(t)}{\|z(t)\|},\frac{d}{dt}z(t)\rangle|\leq\|u_{t}(\psi_{t}(x\mid h)\mid h)-u_{t}(\psi_{t}(x^{\prime}\mid h)\mid h)\|$ (62)
@@
$\frac{d}{dt}d(t)\leq\|u_{t}(\psi_{t}(x\mid h)\mid h)-u_{t}(\psi_{t}(x^{\prime}\mid h)\mid h)\|$ (63)
Since the guided vector field is Lipschitz continuous,
$\frac{d}{dt}d(t)\leq L_{u}d(t)$ (64)
Based on Gronwall’s inequality [gronwall1919note, hirsch2013differential],
Assuming that $d(t)>0$ divide both sides by $d(t)$. If $d(t)=0$, the inequality holds.
$\frac{1}{d(t)}\frac{d}{dt}d(t)\leq L\Rightarrow\frac{d}{dt}\log d(t)\leq L$ (65)
Now integrate both sides from 0 to $t$:
$\log d(t)-\log d(0)\leq Lt\Rightarrow\log\left(\frac{d(t)}{d(0)}\right)\leq Lt\Rightarrow\frac{d(t)}{d(0)}\leq e^{Lt}\Rightarrow d(t)\leq d(0)e^{Lt}$ (66)
Since $d(0)=\|\psi_{0}(x\mid h)-\psi_{0}(x^{\prime}\mid h)\|=\|x-x^{\prime}\|$,
$\|\psi_{t}(x\mid h)-\psi_{t}(x^{\prime}\mid h)\|\leq e^{L_{u}t}\|x-x^{\prime}\|$ (67)
Therefore, we know that
$\|\psi_{t}(x\mid h)-\psi_{t}(x^{\prime}\mid h)\|\leq e^{L_{u}}\|x-x^{\prime}\|\quad\forall x,x^{\prime},t,h$ (68)
@@
where $\det A$ denotes the determinant of a square matrix $A$ and $J_{\psi^{-1}}(y)=\frac{\partial\psi^{-1}(y)}{\partial y}$ is the Jacobian of $\psi^{-1}$. The probability mass of the transformed set $\mathcal{A}^{\prime}=\psi(\mathcal{A})$ is:
@@
Assume that the guided vector field is bi-Lipschitz uniformly in $x$ over $t\in[0,1]$ and $h\in\mathcal{H}$, i.e., there exists $L_{u}$ and $l_{u}$ such that
@@
Proof follows similarly to Lemma 58. The upper Lipschitz bound follows from Lemma 58.
@@
Since $\frac{d}{dt}z(t)=u_{t}(x\mid h)-u_{t}(x^{\prime}\mid h)$ and $\|u_{t}(x\mid h)-u_{t}(x^{\prime}\mid h)\|\geq l_{u}\|x-x^{\prime}\|=l_{u}\|\psi_{t}(x\mid h)-\psi_{t}(x^{\prime}\mid h)\|$, we obtain
@@
Combining with the upper Lipschitz bound, we get
@@
Proof of Lemma 23.
The proof follows the proof of Lemma 1 in xu2023conformal. Under the assumption that $\{e_{i}\}_{i=1}^{T+1}$ are i.i.d., the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality [dvoretzky1956asymptotic, kosorok2008introduction] implies:
@@
Lemma A.10 (Gaussian concentration inequality, Theorem 5.6 in boucheron2003concentration).
@@
$\mathbb{P}(f(X)\geq\mathbb{E}[f(X)]+t)\leq\exp\left(\frac{-t^{2}}{2{L_{f}}^{2}}\right),$ (85)
@@
Let $X\sim\mathcal{N}(0,\gamma I_{d})$ be an isotropic Gaussian random vector in $\mathbb{R}^{d}$ with covariance matrix $\gamma I_{d}\in\mathbb{R}^{d}$ for some $\gamma>0$ and let $f:\mathbb{R}^{d}\to\mathbb{R}$ be an $L_{f}$-Lipschitz continuous function. Then, for all $t>0$,
$\mathbb{P}(f(X)\geq\mathbb{E}[f(X)]+t)\leq\exp\left(\frac{-t^{2}}{2\gamma{L_{f}}^{2}}\right),$ (86)
@@
Let $X^{\prime}\sim\mathcal{N}(0,I_{d})$, and define $X=\sqrt{\gamma}X^{\prime}$, so that $X\sim\mathcal{N}(0,\gamma I_{d})$. Define the function $f_{\gamma}(x):=f(\sqrt{\gamma}x)$. Then $f_{\gamma}$ is $\sqrt{\gamma}L_{f}$-Lipschitz. Applying Lemma A.10 to $f_{\gamma}(X^{\prime})$, we obtain:
$\mathbb{P}\left(f_{\gamma}(X^{\prime})\geq\mathbb{E}[f_{\gamma}(X^{\prime})]+t\right)\leq\exp\left(-\frac{t^{2}}{2\gamma L_{f}^{2}}\right).$ (87)
@@
$\mathbb{P}(f(X)\geq\mathbb{E}[f(X)]+t)=\mathbb{P}\left(f_{\gamma}(X^{\prime})\geq\mathbb{E}[f_{\gamma}(X^{\prime})]+t\right)\leq\exp\left(-\frac{t^{2}}{2\gamma L_{f}^{2}}\right).$ (88)
Proof of Lemma 4.11.
Let $X\sim\mathcal{N}(0,\gamma I_{d})$ be an isotropic Gaussian random vector in $\mathbb{R}^{d}$ with covariance matrix $\gamma I_{d}\in\mathbb{R}^{d}$ for some $\gamma>0$ and let $f:\mathbb{R}^{d}\to\mathbb{R}$ be 2-norm, i.e., $f(X)=\|X\|$.
@@
Under Assumption 4.1, 4.6, and 4.9, with probability at least $1-\delta$,
$\sum_{i=1}^{T}|\hat{e}_{i}-e_{i}|\leq 2T(M_{T}L_{\psi^{-1}}\delta_{T}+L_{\psi^{-1}}^{2}\delta^{2}_{T}).$ (96)
@@
Using the identity for the difference of squared norms:
@@
$=\|s_{i}\|^{2}+2\langle s_{i},\hat{s_{i}}-s_{i}\rangle+\|\hat{s_{i}}-s_{i}\|^{2},$
we obtain:
$\|\hat{s_{i}}\|^{2}-\|s_{i}\|^{2}=2\langle s_{i},\hat{s_{i}}-s_{i}\rangle+\|\hat{s_{i}}-s_{i}\|^{2}$ (98)
@@
By Lemma 4.11, we have with probability at least $1-\delta$ that $\|s_{i}\|\leq M_{T}$ for all $i$, and by Assumption 4.9, $\|\Delta_{i}\|\leq\delta_{T}$. Substituting these into the inequality (103),
@@
Summing over all $i=1,\dots,T$, we conclude:
@@
Proof of Lemma 4.12.
By Lemma A.12, we have with probability at least $1-\delta$,
@@
$\overset{(i)}{=}\left|\mathbb{P}\left(\hat{e}_{T+1}\in\left[\widehat{F}_{T+1}^{-1}(\beta),\,\widehat{F}_{T+1}^{-1}(1-\alpha+\beta)\right]\mid Z_{T+1}=z_{T+1}\right)-(1-\alpha)\right|$
$\overset{(ii)}{=}\left|\mathbb{P}\left(\beta\leq\widehat{F}_{T+1}(\hat{e}_{T+1})\leq 1-\alpha+\beta\right)-\mathbb{P}\left(\beta\leq F_{e}(e_{T+1})\leq 1-\alpha+\beta\right)\right|.$
In equality $(i)$, we consider the transformation of the prediction residual into $\hat{e}_{T+1}$ conditioned on $h_{T+1}=\text{Enc}(z_{T+1})$ approximates $\hat{F}_{T+1}$. Equality $(ii)$ follows from Lemma A.9, which states that $F_{e}(e_{T+1})\sim\mathrm{Unif}[0,1]$. This can be further bounded by:
@@
Here, inequality $(i)$ holds due to the supremum of $|\widehat{F}_{T+1}(x)-F_{e}(x)|$ over $x$ and Lipschitz continuity of $F_{e}$ from Assumption 4.8. Inequality $(ii)$ follows from Lemma 4.12. Inequality $(iii)$ follows from Lemma 23.
@@
Proof of Lemma 28.
The proof follows similarly in the proof of Lemma B.11 in xu2024conformal. Define $v_{T}(x):=\sqrt{T}(\tilde{F}_{T+1}(x)-F_{e}(x))$. By using Proposition 7.1 in rio2017asymptotic, we have:
@@
Under Assumption 4.16, the result follows by combining Lemma 4.12 and 28, using an argument analogous to the proof of Theorem 26.

### v3

@@ 6 Conclusion
In this study, we proposed a novel conformal prediction method for multi-dimensional time series using flow with classifier-free guidance. We provided coverage guarantees of our method by establishing exact non-asymptotic marginal coverage and a finite-sample bound on conditional coverage. Experiments on real-world datasets with a broad set of baselines demonstrated that our method constructs smaller prediction sets while satisfying the target coverage, consistently outperforming the baselines. Future work will focus on investigating flow with optimal transport coupling to construct more efficient prediction sets, and on conducting extensive evaluations across diverse datasets with higher-dimensional outcomes.
@@ Acknowledgments
Reproducibility Statement
The source code for the method and experiments is available at https://github.com/Jayaos/flow_cp. Details of the experimental setup, including hyperparameters, datasets, are provided in the Experiments section and the Appendix B.
@@ Appendix A Proofs
Since $u_{t|x_{1}}(x\mid x_{1})$ generates the probability path $p_{t|x_{1}}(x\mid x_{1})$, the continuity equation holds as follows:
$\frac{\partial}{\partial t}p_{t|x_{1}}(x\mid x_{1})+\text{div}\left(u_{t|x_{1}}(x\mid x_{1})p_{t|x_{1}}(x\mid x_{1})\right)=0.$ (26)
By the definition of the guided marginal probability path, we have:
$p_{t|h}(x\mid h)=\int p_{t|x_{1}}(x\mid x_{1})q(x_{1}\mid h)dx_{1}.$ (27)
The time derivative of $p_{t|h}(x\mid h)$ is expressed as:
$\frac{\partial}{\partial t}p_{t|h}(x\mid h)$ $=\frac{\partial}{\partial t}\int p_{t|x_{1}}(x\mid x_{1})q(x_{1}\mid h)\,dx_{1}$ (28)
$=\int\frac{\partial}{\partial t}p_{t|x_{1}}(x\mid x_{1})q(x_{1}\mid h)\,dx_{1}$
@@
$=-\text{div}\int u_{t|x_{1}}(x\mid x_{1})p_{t|x_{1}}(x\mid x_{1})q(x_{1}\mid h)\,dx_{1}.$
The marginal guided vector field is defined as follows:
$u_{t|h}(x\mid h):=\int u_{t|x_{1}}(x\mid x_{1})\frac{p_{t|x_{1}}(x\mid x_{1})q(x_{1}\mid h)}{p_{t|h}(x\mid h)}\,dx_{1}.$ (29)
We can rewrite the marginal vector field as:
@@
$\frac{\partial}{\partial t}p_{t|h}(x\mid h)=-\text{div}\left(u_{t|h}(x\mid h)p_{t|h}(x\mid h)\right),$ (31)
which is the continuity equation for $p_{t|h}(x\mid h)$ under the guided vector field $u_{t|h}(x\mid h)$. Therefore, $u_{t|h}(x\mid h)$ is a valid vector field generating $p_{t|h}(x\mid h)$. ∎
@@
By the definition of the guided marginal probability path, we have:
@@
where $p_{t|x_{1}}(x\mid x_{1})=\mathcal{N}(x\mid\alpha_{t}x_{1},\sigma_{t}^{2}I)$.
We can express the score function of the guided marginal probability path as:
@@
$\overset{(i)}{=}\int\frac{\nabla_{x}p_{t|x_{1}}(x\mid x_{1})q(x_{1}\mid h)dx_{1}}{p_{t|h}(x\mid h)}$ (35)
$=\int\nabla_{x}\log p_{t|x_{1}}(x\mid x_{1})\frac{p_{t|x_{1}}(x\mid x_{1})q(x_{1}\mid h)}{p_{t|h}(x\mid h)}dx_{1},$ (36)
where $(i)$ follows from the definition of the guided marginal probability path.
@@
$u_{t|x_{1}}(x\mid x_{1})$ $=\frac{\dot{\alpha}_{t}}{\sigma_{t}}(x-\alpha_{t}x_{1})+\dot{\alpha}_{t}x_{1}$ (37)
@@
$=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+(\dot{\alpha}_{t}\sigma_{t}-\alpha_{t}\dot{\sigma}_{t})\frac{\sigma_{t}}{\alpha_{t}}\nabla_{x}\log p_{t|x_{1}}(x\mid x_{1}),$ (41)
@@
Therefore, by pugging $u_{t|x_{1}}(x\mid x_{1})$ into equation (42), we have:
@@
By using the identity $\nabla_{x}\log p_{t|h}(x\mid h)=\nabla_{x}\log p_{h|t}(h\mid x)+\nabla_{x}\log p_{t}(x)$, we obtain:
@@
By using Jacobi’s formula, we have:
@@
$\text{tr}\left(\frac{\partial u_{t|h}(\psi_{t|h}(x\mid h))}{\partial\psi_{t|h}(x\mid h)}\right)=\text{div}\left(u_{t|h}(\psi_{t|h}(x\mid h))\right).$ (50)
@@
Theorem A.4 (Continuous images of connected sets, Rudin [37]).
Let $X$ and $Y$ be topological spaces and $\psi:X\to Y$ be continuous. If $E\subset X$ is connected, then $\psi(E)$ is connected.
Theorem A.5 (Closed sets under homeomorphism, Munkres [33]).
Let $X$ and $Y$ be topological spaces and $\psi:X\to Y$ be a homeomorphism. If $E\subset X$ is closed, then $\psi(E)\subset Y$ is closed.
@@
Let $\psi_{t}$ denote the guided flow defined by a guided vector field $u_{t}$. If the guided vector field $u_{t}(x\mid h)$ is Lipschitz continuous in $x$ for all $t$ and $h$, i.e., there exists a constant $L_{u}>0$ such that
@@
then the guided flow $\psi_{t}(x\mid h)$ is Lipschitz continuous in $x$ for all $t$ and $h$. That is, there exists a constant $L_{\psi}>0$ such that
@@
Let $d(t)=\|\psi_{t}(x\mid h)-\psi_{t}(x^{\prime}\mid h)\|$ and $z(t)=\psi_{t}(x\mid h)-\psi_{t}(x^{\prime}\mid h)$. Since the guided vector field is Lipschitz continuous, there exists $L_{u}$ such that
@@
Since $d(t)\neq 0$, we have:
$\frac{d}{dt}d(t)=\frac{1}{\|z(t)\|}\langle z(t),\frac{d}{dt}z(t)\rangle=\langle\frac{z(t)}{\|z(t)\|},\frac{d}{dt}z(t)\rangle.$ (57)
Since $\frac{d}{dt}z(t)=u_{t}(\psi_{t}(x\mid h)\mid h)-u_{t}(\psi_{t}(x^{\prime}\mid h)\mid h)$, by Cauchy-Schwarz inequality, we have:
$|\langle\frac{z(t)}{\|z(t)\|},\frac{d}{dt}z(t)\rangle|\leq\|u_{t}(\psi_{t}(x\mid h)\mid h)-u_{t}(\psi_{t}(x^{\prime}\mid h)\mid h)\|,$ (58)
which implies:
$\frac{d}{dt}d(t)\leq\|u_{t}(\psi_{t}(x\mid h)\mid h)-u_{t}(\psi_{t}(x^{\prime}\mid h)\mid h)\|\leq L_{u}d(t),$ (59)
where the last inequality follows from the Lipschitz continuity of the guided vector field.
Based on Gronwall’s inequality [17, 20], it follows that
$d(t)\leq d(0)\,e^{L_{u}t}.$ (60)
This can be proved by defining $\phi(t):=e^{-L_{u}t}d(t)$. Then,
$\frac{d}{dt}\phi(t)=e^{-L_{u}t}\big(\frac{d}{dt}d(t)-L_{u}d(t)\big)\leq 0,$ (61)
which implies $\phi$ is non-increasing on $t$. Since $\phi(t)\leq\phi(0)=d(0)$ for all $t\in[0,1]$, multiplying both sides by $e^{L_{u}t}$ yields:
$d(t)\leq d(0)e^{L_{u}t}.$ (62)
@@
$\log d(t)-\log d(0)\leq L_{u}t.$ (63)
Exponentiating both sides gives
$\frac{d(t)}{d(0)}\leq e^{L_{u}t},$ (64)
and hence
$d(t)\leq d(0)e^{L_{u}t}.$ (65)
Since $d(0)=\|\psi_{0}(x\mid h)-\psi_{0}(x^{\prime}\mid h)\|=\|x-x^{\prime}\|$, we conclude that
$\|\psi_{t}(x\mid h)-\psi_{t}(x^{\prime}\mid h)\|\leq e^{L_{u}t}\|x-x^{\prime}\|.$ (66)
@@
where $\det A$ denotes the determinant $A$ and $J_{\psi^{-1}}(y)=\frac{\partial\psi^{-1}(y)}{\partial y}$ is the Jacobian of $\psi^{-1}$. The probability mass of the transformed set $\mathcal{A}^{\prime}=\psi(\mathcal{A})$ is:
@@
Assume that the guided vector field is bi-Lipschitz in $x$ over for all $t$ and $h$, i.e., there exists $L_{u}$ and $l_{u}$ such that
@@
Proof follows similarly to Lemma 54.
@@
Since $\frac{d}{dt}z(t)=u_{t}(x\mid h)-u_{t}(x^{\prime}\mid h)$ and $\|u_{t}(x\mid h)-u_{t}(x^{\prime}\mid h)\|\geq l_{u}\|x-x^{\prime}\|=l_{u}\|\psi_{t}(x\mid h)-\psi_{t}(x^{\prime}\mid h)\|$, we have:
@@
Combining with the upper Lipschitz bound obtained from Lemma 54, we conclude that
@@
Lemma A.9 (Convergence of empirical CDF of i.i.d. $\{e_{i}\}_{i=1}^{T}$).
Proof.
This proof follows the proof of Lemma 1 in Xu and Xie [53]. Under the assumption that $\{e_{i}\}_{i=1}^{T+1}$ are i.i.d., the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality [14, 27] implies:
@@
Lemma A.10 (Gaussian concentration inequality, Theorem 5.6 in Boucheron et al. [6]).
@@
$\mathbb{P}(f(X)\geq\mathbb{E}[f(X)]+t)\leq\exp\left(\frac{-t^{2}}{2{L_{f}}^{2}}\right).$ (84)
@@
Let $X\sim\mathcal{N}(0,\gamma I_{d})$ be an isotropic Gaussian random vector in $\mathbb{R}^{d}$ with covariance matrix $\gamma I_{d}\in\mathbb{R}^{d}$ for some $\gamma>0$. Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be an $L_{f}$-Lipschitz continuous function. Then, for all $t>0$,
$\mathbb{P}(f(X)\geq\mathbb{E}[f(X)]+t)\leq\exp\left(\frac{-t^{2}}{2\gamma{L_{f}}^{2}}\right).$ (85)
@@
Let $X^{\prime}\sim\mathcal{N}(0,I_{d})$, and define $X=\sqrt{\gamma}X^{\prime}$, so that $X\sim\mathcal{N}(0,\gamma I_{d})$. Define the function $f_{\gamma}(x):=f(\sqrt{\gamma}x)$. Then $f_{\gamma}$ is $\sqrt{\gamma}L_{f}$-Lipschitz. Using Lemma A.10 to $f_{\gamma}(X^{\prime})$, we obtain:
$\mathbb{P}\left(f_{\gamma}(X^{\prime})\geq\mathbb{E}[f_{\gamma}(X^{\prime})]+t\right)\leq\exp\left(\frac{-t^{2}}{2\gamma{L_{f}}^{2}}\right).$ (86)
@@
$\mathbb{P}(f(X)\geq\mathbb{E}[f(X)]+t)=\mathbb{P}\left(f_{\gamma}(X^{\prime})\geq\mathbb{E}[f_{\gamma}(X^{\prime})]+t\right)\leq\exp\left(\frac{-t^{2}}{2\gamma{L_{f}}^{2}}\right).$ (87)
Lemma A.12 (Norm concentration of isotropic Gaussian random vectors).
$\max_{1\leq i\leq T}\|X_{i}\|\leq M_{T},$ (88)
Proof.
@@
Under Assumption 4.6 and 4.9, with probability at least $1-\delta$,
$\sum_{i=1}^{T}|\hat{e}_{i}-e_{i}|\leq T(2M_{T}L_{\psi^{-1}}\delta_{T}+L_{\psi^{-1}}^{2}\delta^{2}_{T}).$ (96)
@@
Using the identity for the difference of squared norms, we have:
@@
$=\|s_{i}\|^{2}+2\langle s_{i},\hat{s_{i}}-s_{i}\rangle+\|\hat{s_{i}}-s_{i}\|^{2}.$
@@
By Lemma A.12, we have with probability at least $1-\delta$ that $\|s_{i}\|\leq M_{T}$ for all $i$, and by Assumption 4.9, $\|\Delta_{i}\|\leq\delta_{T}$. Substituting these into the inequality (103), we have:
@@
Summing over all $i=1,\dots,T$, we conclude that
@@
Lemma A.14 (Distance between the empirical CDF of $\{e_{i}\}_{i=1}^{T}$ and $\{\hat{e}_{i}\}_{i=1}^{T}$).
where $C=\sqrt{M_{T}L_{\psi^{-1}}\delta_{T}+L_{\psi^{-1}}^{2}\delta_{T}^{2}}$.
Proof.
By Lemma A.13, with probability at least $1-\delta$, we have:
@@
$=\left|\mathbb{P}\left(\hat{e}_{T+1}\in\left[\widehat{F}_{T+1}^{-1}(\beta),\,\widehat{F}_{T+1}^{-1}(1-\alpha+\beta)\right]\mid Z_{T+1}=z_{T+1}\right)-(1-\alpha)\right|$
$\overset{(i)}{=}\left|\mathbb{P}\left(\beta\leq\widehat{F}_{T+1}(\hat{e}_{T+1})\leq 1-\alpha+\beta\right)-\mathbb{P}\left(\beta\leq F_{e}(e_{T+1})\leq 1-\alpha+\beta\right)\right|.$
Equality $(i)$ follows from Lemma A.8, which states that $F_{e}(e_{T+1})\sim\mathrm{Unif}[0,1]$. This can be further bounded by:
@@
Here, inequality $(i)$ holds due to the supremum of $|\widehat{F}_{T+1}(x)-F_{e}(x)|$ over $x$ and Lipschitz continuity of $F_{e}$ from Assumption 4.8. Inequality $(ii)$ follows from Lemma A.14. Inequality $(iii)$ follows from Lemma 80.
@@
Definition A.15.
Definition A.16.
Lemma A.17 (Convergence of empirical CDF of stationary and strongly mixing $\{e_{i}\}_{i=1}^{T}$).
Proof.
The proof follows similarly in the proof of Lemma B.11 in Xu et al. [50]. Define $v_{T}(x):=\sqrt{T}(\tilde{F}_{T+1}(x)-F_{e}(x))$. By using Proposition 7.1 in Rio and others [36], we have:
@@
Under Assumption 4.11, the result follows by combining Lemma A.14 and 123, using an argument analogous to the proof of Theorem 23.